OpenStudy (jenniferjuice):

Choose the correct simplification of the expression (5x^4y^2z^2)(3x^4y^3z^5).

OpenStudy (uri):

(5x^4y^2z^2)*(3x^4y^3z^5). What is 5*3?

OpenStudy (jenniferjuice):

15

OpenStudy (anonymous):

are they all multiplied by each other?

OpenStudy (anonymous):

in the parenthesis?

OpenStudy (uri):

Now add all the powers x^4+x^4=x^8 So what is y^2+y^3=? and z^2+z^5=?

OpenStudy (uri):

It will be 15x^8y^5z7